Sheyanne17:

A garden store has the following miscellaneous flower bulbs in a basket 5 amaryllis 6 daffodils 4 lilies 3 tulips A customer bought 4 bulbs from the basket, one of each type of flower. If the next customer selects 1 of the remaining bulbs at random, which is closest to the probability that customer will get a daffodil bulb? A.) 36% B.) 27% C.) 29% D.) 21%

celticcat:

after the first purchase there 4 amaryllis , 5 daffodils 3 lilies and 2 tulips A total of 14 So Prob ( next customer picks a daffodil) = 5 / 14 = 36% to nearest percent

FruitDuit:

@celticcat I believe you are correct. I will simply write it out in more detail. One person has bought one of each bulb in the garden store. By the time a second person comes in to buy a bulb, you will have 4 amaryllis, 5 daffodils, 3 lilies, and 2 tulips. The probability of obtaining a certain bulb is the number of bulbs (of that species) divided by the total number of different bulbs. This will be 5 (daffodils)/14 (total number of bulbs). 5/14 = 0.3571 and so on. To convert a decimal to a percent, multiply it by 100 OR simply move the decimal over two places (to the right). You now have 35.71%. Due to the seven, you can round the percent UP to 36% and obtain your answer, which is A.